OpenStudy (anonymous):

I have been confused on this for a while... 2x-3y=-1 y=x-1 Please help ;-;

jhonyy9 (jhonyy9):

@Masay so this is very easy hope that you need to solve this system for x and y - so than just substitute the y=x-1 in the first equation in place of y and will get the x so after that just substitute this value of x in place of x inside second equation and will get the value of y - these are all hope this will help you to solve your exercise easy

jhonyy9 (jhonyy9):

do you understand it what i have wrote above ?

jhonyy9 (jhonyy9):

are you here ?

jhonyy9 (jhonyy9):

courage please - i will help you step by step

jhonyy9 (jhonyy9):

2x-3y=-1 y=x-1 so i have wrote above just you need substitute the second equation y=x-1 inside first equation in place of y - so what will get ?

OpenStudy (anonymous):

I am , I was finishing the problem, and apparently the reason I was confused was due to how I accidentally didn't add a subtraction sign! Whoops! Thank you for helping ^^ also in place of y you would put x-1

OpenStudy (anonymous):

which would be 2x-3 ( x-1 )= -1

jhonyy9 (jhonyy9):

yes and continue it - or you have finished it right ?

jhonyy9 (jhonyy9):

ok. so than you are welcome was my pleasure -- good luck - bye

OpenStudy (anonymous):

Thank you~ ^^